### v1

@@ 2.4 Training a single-lead ECG model based on leads augmentation
The development of portable and wearable ECG devices in recent years has revolutionized continuous cardiac monitoring, offering a non-invasive method for real-time health assessment. Beyond the conventional diagnosis of cardiac arrhythmias, another critical challenge in this field is accurately detecting and interpreting ECG axis deviation on single-lead ECGs from portable devices (typically lead I), which can significantly impact the diagnosis of various cardiac abnormalities. For instance, left axis deviation can provide additional insights into diagnoses like left ventricular hypertrophy, left bundle branch block, left anterior fascicular block, pre-excitation syndromes, and inferior myocardial infarction.[17, 8] To address this issue, we developed a novel training method utilizing lead-augmented portable ECG models. By systematically enhancing standard 12-lead ECG data, we simulate various clinical scenarios of axis inversion, thereby enhancing the model’s robustness and versatility. Understanding the relationship between ECG vectors and leads is crucial for this method. The cardiac electrical activity generates a vector representation of cardiac signals, captured by different ECG leads placed on the body. Each lead provides a unique perspective of the cardiac electrical axis, offering a comprehensive view when combined. The standard 12-lead ECG system includes limb leads (I, II, III, aVR, aVL, aVF) and precordial leads (V1-V6), with each lead representing a specific projection of the cardiac electrical vector.[22] Specifically, we primarily utilize the ECG signals from limb leads. Based on the angular position of each limb lead’s axis relative to the heart, we consider lead I as the center of the semicircle (i.e., 0°) and calculate the signals from all leads around the semicircle (i.e., from -90° to 90°), thereby obtaining six additional leads (aVL, -aVR, II, -III, aVF, -aVF). We then trained a model for portable ECG devices, extracting the lead I ECG from the HEEDB 12-lead data and randomly incorporating one of the remaining six augmented leads into the model for training with a 50% probability. This ensures that the model can learn arrhythmia features from lead I data and axis deviation from the additional six leads. Additionally, we have scaled down the model’s parameter size to optimize for portable devices with limited computational resources.
@@ 2.5 Fine-tuning on ECG foundation model
When adapting to specific ECG downstream tasks, we need to retain the parameters of the base model and discard the initial classification linear layer. The number of classes in the downstream task determines the number of neurons needed in the final layer of the new linear layer. The training objective is to generate classification outputs that match the labels. To fully leverage the potential of the foundation model, we experimented with two fine-tuning strategies. The first fine-tuning strategy involved loading the pre-trained model parameters and fine-tuning the entire model. The second fine-tuning strategy involved loading the model parameters, freezing the encoder’s parameters (i.e., not altering the encoder parameters), and only fine-tuning the final linear layer of the model.
The batch size is 256. The total training period is 30 epochs, with a learning rate adjustment strategy that utilizes the scheduler. After every epoch, the scheduler monitors the specified metric, and if the performance does not improve for 10 consecutive epochs, the learning rate is reduced by a factor of 0.1. The learning rate reduction is triggered based on the maximization of the monitored metric. This approach ensures a dynamic adjustment of the learning rate depending on the training progress. After training in each epoch, the model is evaluated on the validation set.[4] The model weights with the highest AUC on the validation set are saved as a checkpoint for future evaluation.
@@ 2.6 Clinical validation
2.7 Statistical analysis
The evaluation of the model was conducted by calculating accuracy, the area under the Receiver Operating Characteristic (ROC) curve (AUC), sensitivity, specificity, and the F1 score, each with bilateral 95% confidence intervals (CI).
@@ 3 Results
Our dataset comprises 10,771,552 expert-annotated ECGs recorded from 1,818,247 unique subjects using 12-lead ECG machines, covering the period from the 1990s to present, with a recording duration of 10 seconds. The average age of patients is 60.6 years, with 48.7% being female.[19] After annotation cleaning, the training dataset became a multi-label dataset, with over 10,558,162 (98%) of ECGs being multi-labeled. The development dataset includes 7,519,035 ECGs from 1,725,643 patients, and the test dataset includes 834,926 ECGs from 265,818 patients. The committee internal test set includes 523 ECGs recorded between January 10, 2023, and January 20, 2023, corresponding to 523 unique patients. We obtained a total of 150 labels describing the ECGs, including heart rate abnormalities, conduction abnormalities, cardiac hypertrophy, myocardial infarction, myocardial ischemia, pacemaker devices, and ECG waveform morphology. Using these labels, we trained two deep learning ECG models separately: one for standard 12-lead ECGs and the other designed for single-lead portable device ECGs (lead I). We evaluated the performance of both models on internal and external test sets.
@@
We validated the deep learning model for 12-lead ECGs using the committee internal test set, which includes 523 ECGs recorded between January 10, 2023, and January 20, 2023, corresponding to 523 unique patients (49% female; mean age 57.2 years). In internal test set experiments for 12-lead ECGs, the algorithm’s average AUC score for diagnosing all 20 classifications was 0.968 (95% CI, 0.965–0.969), the PRAUC score was 0.912 (95% CI, 0.910–0.914), sensitivity was 0.971 (95% CI, 0.970–0.973), and specificity was 0.937 (95% CI, 0.935–0.938). When comparing the model with cardiologists, our model achieved an overall average F1 score of 0.683, outperforming the cardiologists’ average F1 score of 0.640. The model’s performance was compared with cardiologists’ performance across 20 diagnostic categories (Table LABEL:tab:internal). The model’s F1 scores were higher than those of the cardiologists in 12 categories (atrial fibrillation, atrial flutter, incomplete right bundle branch block, lateral infarct, left bundle branch block, normal sinus rhythm, premature atrial complexes, right axis deviation, right bundle branch block, sinus bradycardia, sinus rhythm, with sinus arrhythmia). Notably, the model’s F1 scores for atrial flutter, incomplete right bundle branch block, lateral infarct, left bundle branch block, and right axis deviation were significantly higher than those of the cardiologists. The model’s F1 scores for anterior infarct, atrial-paced rhythm, inferior infarct, and ventricular tachycardia were slightly lower than those of the cardiologists. When comparing the model’s ROC curve with the true positive rate and false positive rate of cardiologists, the model outperformed the average performance of cardiologists for most labels.
@@
In the external test set experiments on single-lead ECG devices, our model accurately classified sinus rhythm and atrial fibrillation. The model’s performance, as shown in the table, achieved an AUC of 0.975 and 0.957 for normal sinus rhythm and atrial fibrillation, respectively (Table 10). This indicates that the model achieves excellent performance on portable and wearable ECGs in real-world scenarios.
@@
We next validated the performance of ECGFounder in transfer learning. Our model was fine-tuned and evaluated on six downstream tasks using supervised learning on the MIMIC-IV-ECG dataset, resulting in both a 12-lead ECG model and a single-lead ECG model. The fine-tuning results are shown in Figures 3. As shown, ECGFounder outperforms the baseline methods in every downstream task. Specifically, it achieves 2 to 3 percentage points higher performance than ECG-SimCLR and 4 to 6 points higher than ECG-ResNet in age, sex, NTproBNP, CKD, and CHD detection. Additionally, to compare with the previously published CKD study [13], we created an independent test set consisting of individual patients from the MIMIC-IV-ECG dataset for validation. As shown in the results (Table 2), our model surpasses the previously studied model in all six classifications for both 12-lead and single-lead ECGs.
@@
Table 1: Performance of other ECG deep learning model and ECGFounder on external test set CODE-test and PTB-XL.[31, 38] In particular, for the baseline method S12L-ECG, the CODE-test is the internal test set.
@@
Dataset CODE-test Models ECGFounder S12L-ECG(Internal)[31] CTN[24] ECG-SE-ResNet[43] Count AUC Sens Spec AUC Sens Spec AUC Sens Spec AUC Sens Spec SINUS BRADYCARDIA 16 0.967 1.000 0.955 0.955 0.938 0.996 0.965 0.987 0.942 0.932 0.995 0.937 ATRIAL FIBRILLATION 13 0.999 1.000 0.996 0.963 0.769 1.000 0.966 0.972 0.969 0.976 0.980 0.970 SINUS TACHYCARDIA 37 0.989 0.974 0.970 0.977 0.973 0.997 0.972 0.958 0.943 0.976 0.950 0.957 RIGHT BUNDLE BRANCH BLOCK 34 0.989 0.971 0.971 0.995 1.000 0.995 0.989 0.946 0.986 0.965 0.949 0.947 LEFT BUNDLE BRANCH BLOCK 30 0.998 1.000 0.996 1.000 1.000 1.000 0.961 0.988 0.942 0.971 0.949 0.976 WITH 1ST DEGREE AV BLOCK 28 0.949 0.864 0.957 0.989 0.929 0.995 0.930 0.862 0.925 0.958 0.856 0.922 Dataset PTB-XL Models ECGFounder S12L-ECG[31] CTN[24] ECG-SE-ResNet[43] ANTERIOR INFARCT 36 0.635 0.816 0.447 / / / / / / / / / ANTEROLATERAL INFARCT 29 0.945 0.837 0.918 / / / / / / / / / ANTEROSEPTAL INFARCT 2415 0.945 0.891 0.882 / / / / / / / / / ATRIAL FIBRILLATION 1514 0.993 0.969 0.982 0.925 0.776 0.989 0.953 0.935 0.915 0.950 0.946 0.907 ATRIAL FLUTTER 73 0.993 0.973 0.969 / / / 0.948 0.962 0.937 0.941 0.966 0.956 ELECTRONIC ATRIAL PACEMAKER 298 0.909 0.935 0.745 / / / / / / / / / INFERIOR INFARCT 276 0.851 0.825 0.732 / / / / / / / / / LATERAL INFARCT 1066 0.916 0.819 0.859 / / / / / / / / / LEFT ANTERIOR FASCICULAR BLOCK 167 0.972 0.922 0.919 / / / 0.965 0.882 0.914 0.968 0.819 0.901 LEFT ATRIAL ENLARGEMENT 427 0.799 0.742 0.704 / / / / / / / / / LEFT BUNDLE BRANCH BLOCK 539 0.989 0.969 0.965 0.983 0.963 0.991 0.912 0.932 0.933 0.931 0.917 0.930 LEFT POSTERIOR FASCICULAR BLOCK 18 0.840 0.605 0.904 / / / / / / / / / LEFT VENTRICULAR HYPERTROPHY 2419 0.900 0.860 0.788 / / / / / / / / / LOW VOLTAGE QRS 182 0.581 0.690 0.423 / / / 0.605 0.605 0.814 0.593 0.655 0.741 N-SPECIFIC INTRAVENTRICULAR BLOCK 78 0.766 0.339 0.932 / / / 0.788 0.487 0.906 0.776 0.699 0.793 NORMAL ECG 987 0.887 0.952 0.639 / / / / / / / / / PREMATURE VENTRICULAR COMPLEXES 1146 0.987 0.965 0.959 / / / 0.963 0.955 0.921 0.937 0.939 0.880 QT HAS LENGTHENED 119 0.931 0.897 0.827 / / / 0.930 0.881 0.957 0.930 0.901 0.930 RIGHT ATRIAL ENLARGEMENT 99 0.959 0.869 0.915 / / / / / / / / / RIGHT BUNDLE BRANCH BLOCK 1155 0.976 0.925 0.925 0.967 0.911 0.976 0.932 0.922 0.931 0.946 0.842 0.912 RIGHT VENTRICULAR HYPERTROPHY 136 0.913 0.833 0.859 / / / / / / / / / SEPTAL INFARCT 1423 0.947 0.894 0.880 / / / / / / / / / SINUS BRADYCARDIA 119 0.950 0.972 0.955 0.967 0.938 0.989 0.996 0.966 0.946 0.983 0.935 0.963 SINUS RHYTHM 16785 0.923 0.936 0.751 / / / 0.977 0.942 0.920 0.963 0.926 0.913 SINUS TACHYCARDIA 826 0.994 0.976 0.987 0.988 0.946 0.996 0.986 0.957 0.968 0.979 0.932 0.944 SUPRAVENTRICULAR TACHYCARDIA 27 0.995 0.976 0.959 / / / / / / / / / VENTRICULAR TACHYCARDIA 24 0.987 0.905 0.993 / / / / / / / / / WITH 1ST DEGREE AV BLOCK 82 0.911 0.661 0.921 0.962 0.925 0.992 0.943 0.783 0.911 0.930 0.735 0.893 WITH QRS WIDENING 45 0.618 0.414 0.742 / / / / / / / / / WOLFF-PARKINSON-WHITE 18 0.919 0.747 0.982 / / / / / / / / /
@@
Models ECGFounder Ouyang[13]
@@ 4 Discussion
ECGFounder enhances diagnostic performance by learning to identify ECG features associated with cardiovascular diseases, which are typically diagnosed based on specific waveform patterns and rhythm characteristics like the elevated ST segments of myocardial infarction and the irregular fluctuations of atrial fibrillation. These features involve anomalies in cardiac electrical activity, appearing significantly different from normal ECG waveforms. Upon training, ECGFounder can recognize these disease-related waveform patterns and rhythms, accurately diagnosing corresponding cardiovascular conditions. As observed in Tables LABEL:tab:internal and Figure 2, ECGFounder matches or even exceeds the performance of cardiology experts on the internal review test sets. Furthermore, ECGFounder’s sensitivity exceeds that of the average cardiology expert, indicating its ability to more accurately capture subtle signs of cardiovascular diseases that may be overlooked by human experts.
@@
Previous research has demonstrated that deep learning models can support clinical ECG analysis and achieve good performance.[11] However, existing models lack a universal clinical diagnostic capability for ECGs. Firstly, the training datasets used by existing models are not large or diverse enough, which can lead to overfitting and poor performance on new data, thus limiting their generalizability.[33] Additionally, the lack of demographic diversity in the training datasets means these models may perform poorly for certain demographic groups. This could lead to biases when the models are applied to data from different demographic backgrounds, affecting the fairness and accuracy of the models. Secondly, the labels for most model training datasets are derived from manual annotations by cardiology experts, which is time-consuming and labor-intensive. This limits the number of ECGs available for training. Also, because cardiology experts typically use a unified annotation system, the richness of the dataset labels is not very high, often only including common ECG abnormalities and omitting many important but rare diagnostic labels.[35] Thirdly, due to the training methods used, these models do not support both 12-lead and single-lead ECGs. To address these issues, we propose a foundation model for ECGs that supports both 12-lead and single-lead ECGs. In Tables 3 and 4, we observed that ECGFounder ranks first in average performance across various external tests. The other baseline models, including S12L-ECG, CTN, and ECG-SE-ResNet, have previously achieved the best performance in AI-ECG models.[24, 43, 31] S12L-ECG was trained using supervised learning on the CODE-ECG dataset, which includes 2 million ECGs with 6 common types of arrhythmia diagnoses.[31] CTN and ECG-SE-ResNet were trained on the PhysioNet Challenge-2020 dataset, which includes 60,000 ECGs covering 27 common ECG diagnoses.[43, 24] In this paper, we demonstrate that by training on a larger and more diverse ECG dataset, a scalable foundation model can be developed to further improve the diagnosis of cardiovascular diseases and surpass previous baseline models.
@@
Despite the effectiveness of ECGFounder in detecting various cardiovascular diseases, there are still challenges to be addressed. Firstly, as most data used to develop ECGFounder come from U.S. cohorts, this limits the diversity and representativeness of the data. Different regions of the world may have unique ECG characteristics, such as race and regional-specific rhythm variations, requiring the model to handle data from diverse backgrounds and populations. Secondly, although AI models can provide high-accuracy diagnostic results, their decision-making process is often a ’black box,’ which might encounter trust and acceptance issues in clinical applications. Therefore, developing explainable AI models, allowing doctors to understand the model’s decision-making process, is key to advancing the use of ECG AI models. Lastly, some clinically relevant information, such as patients’ lifestyles and medical histories, which could serve as effective covariates in cardiovascular disease research, have not yet been included in the model. We suggest that future work should involve incorporating a larger volume of ECG data from different regions and ethnicities, adding demographic information of patients as model inputs, and developing more explainable AI models, enabling doctors to better understand the decision processes and outcomes of the models. At the same time, some of the latest natural language processing methods have shown great potential in handling ECG annotations from experts, allowing for a more effective use of the clinical knowledge embedded in these annotations for learning.[20]

### v2

@@ 2.4 Training a single-lead ECG model based on leads augmentation
The development of portable and wearable ECG devices in recent years has revolutionized continuous cardiac monitoring, offering a non-invasive method for real-time health assessment. Beyond the conventional diagnosis of cardiac arrhythmias, another critical challenge in this field is accurately detecting and interpreting ECG axis deviation on single-lead ECGs from wearable devices (typically lead I), which can significantly impact the diagnosis of various cardiac abnormalities. For instance, left axis deviation can provide additional insights into diagnoses like left ventricular hypertrophy, left bundle branch block, left anterior fascicular block, pre-excitation syndromes, and inferior myocardial infarction.27 To address this issue, we developed a novel training method utilizing lead-augmented wearable ECG models. By systematically enhancing standard 12-lead ECG data, we simulate various clinical scenarios of axis inversion, thereby enhancing the model’s robustness and versatility. Understanding the relationship between ECG vectors and leads is crucial for this method. The cardiac electrical activity generates a vector representation of cardiac signals, captured by different ECG leads placed on the body. Each lead provides a unique perspective of the cardiac electrical axis, offering a comprehensive view when combined. The standard 12-lead ECG system includes limb leads (I, II, III, aVR, aVL, aVF) and precordial leads (V1-V6), with each lead representing a specific projection of the cardiac electrical vector.28 Specifically, we primarily utilize the ECG signals from limb leads. Based on the angular position of each limb lead’s axis relative to the heart, we consider lead I as the center of the semicircle (i.e., 0°) and calculate the signals from all leads around the semicircle (i.e., from -90° to 90°), thereby obtaining six additional leads (aVL, -aVR, II, -III, aVF, -aVF). We then trained a model for wearable ECG devices, extracting the lead I ECG from the HEEDB 12-lead data and randomly incorporating one of the remaining six augmented leads into the model for training with a 50% probability. This ensures that the model can learn arrhythmia features from lead I data and axis deviation from the additional six leads. Additionally, we have scaled down the model’s parameter size to optimize for wearable devices with limited computational resources.
@@ 2.5 Fine-tuning on ECG foundation model
When adapting to specific ECG downstream tasks, we need to retain the parameters of the base model and discard the initial classification linear layer. The number of classes in the downstream task determines the number of neurons needed in the final layer of the new linear layer. The training objective is to generate classification outputs that match the labels.
The total training period is 30 epochs, with a learning rate adjustment strategy that utilizes the scheduler. After every epoch, the scheduler monitors the specified metric, and if the performance does not improve for 10 consecutive epochs, the learning rate is reduced by a factor of 0.1. The learning rate reduction is triggered based on the maximization of the monitored metric. This approach ensures a dynamic adjustment of the learning rate depending on the training progress. After training in each epoch, the model is evaluated on the validation set.8 The model weights with the highest AUC on the validation set are saved as a checkpoint for future evaluation.
@@ 2.6 Clinical validation
The evaluation of the model was conducted by calculating accuracy, the area under the Receiver Operating Characteristic curve (AUROC), sensitivity, and specificity, each with bilateral 95% confidence intervals (CI).
@@ 3 Results
Our dataset comprises 10,771,552 expert-annotated ECGs recorded from 1,818,247 unique subjects using 12-lead ECG machines, covering the period from the 1990s to present, with a recording duration of 10 seconds. The average age of patients is 60.6 years, with 48.7% being female.15 The development dataset includes 7,519,035 ECGs from 1,725,643 patients, and the test dataset includes 834,926 ECGs from 265,818 patients. The committee internal test set includes 523 ECGs recorded from 2023, corresponding to 523 unique patients. We obtained a total of 150 labels describing the ECGs, including heart rate abnormalities, conduction abnormalities, cardiac hypertrophy, myocardial infarction, myocardial ischemia, pacemaker devices, and ECG waveform morphology. Using these labels, we trained two deep learning ECG models separately: one for standard 12-lead ECGs and the other designed for single-lead wearable device ECGs (lead I). We evaluated the performance of both models on internal and external test sets.
@@
We validated the deep learning model for 12-lead ECGs using the committee internal test set. The algorithm’s average AUC score for diagnosing all 20 classifications was 0.968 (95% CI, 0.965–0.969), sensitivity was 0.971 (95% CI, 0.970–0.973), and specificity was 0.937 (95% CI, 0.935–0.938). When comparing the model with cardiologists, our model achieved an overall average F1 score of 0.677, outperforming the cardiologists’ average F1 score of 0.640. The model’s performance was compared with cardiologists’ performance across 20 diagnostic categories (Table S7 in Supplement). When comparing the model’s ROC curve with the true positive rate and false positive rate of cardiologists, the model outperformed the average performance of cardiologists for most labels (Figure S2 in Supplement).
@@
In the external test set experiments on single-lead ECG devices, our model accurately classified sinus rhythm and atrial fibrillation. The model’s performance, as shown in the table, achieved an AUC of 0.975 and 0.957 for normal sinus rhythm and atrial fibrillation, respectively (Table S8 in Supplement). This indicates that the model achieves excellent performance on portable and wearable ECGs in real-world scenarios.
@@
We next validated the performance of ECGFounder in transfer learning. Our model was fine-tuned and evaluated on six downstream tasks using supervised learning on the MIMIC-IV-ECG dataset, resulting in both a 12-lead ECG model and a single-lead ECG model. The fine-tuning results are shown in Figures 3. As shown, ECGFounder outperforms the baseline methods in every downstream task. Specifically, it achieves 2 to 3 percentage points higher performance than ECG-SimCLR and 4 to 6 points higher than ECG-ResNet in age, sex, NTproBNP, CKD, and CHD detection. Additionally, to compare with the previously published CKD study 29, we created an independent test set consisting of individual patients from the MIMIC-IV-ECG dataset for validation. As shown in the results (Table 2), our model surpasses the previously studied model in all six classifications for both 12-lead and single-lead ECGs.
@@
Table 1: Performance of other ECG deep learning model and ECGFounder on external test set CODE-test and PTB-XL.17, 18 In particular, for the baseline method S12L-ECG, the CODE-test is the internal test set.
@@
Dataset CODE-test Models ECGFounder S12L-ECG(Internal)17 CTN30 ECG-SE-ResNet31 Count AUC Sens Spec AUC Sens Spec AUC Sens Spec AUC Sens Spec SINUS BRADYCARDIA 16 0.967 1.000 0.955 0.955 0.938 0.996 0.965 0.987 0.942 0.932 0.995 0.937 ATRIAL FIBRILLATION 13 0.999 1.000 0.996 0.963 0.769 1.000 0.966 0.972 0.969 0.976 0.980 0.970 SINUS TACHYCARDIA 37 0.989 0.974 0.970 0.977 0.973 0.997 0.972 0.958 0.943 0.976 0.950 0.957 RIGHT BUNDLE BRANCH BLOCK 34 0.989 0.971 0.971 0.995 1.000 0.995 0.989 0.946 0.986 0.965 0.949 0.947 LEFT BUNDLE BRANCH BLOCK 30 0.998 1.000 0.996 1.000 1.000 1.000 0.961 0.988 0.942 0.971 0.949 0.976 WITH 1ST DEGREE AV BLOCK 28 0.949 0.864 0.957 0.989 0.929 0.995 0.930 0.862 0.925 0.958 0.856 0.922 Dataset PTB-XL Models ECGFounder S12L-ECG17 CTN30 ECG-SE-ResNet31 ANTERIOR INFARCT 36 0.635 0.816 0.447 / / / / / / / / / ANTEROLATERAL INFARCT 29 0.945 0.837 0.918 / / / / / / / / / ANTEROSEPTAL INFARCT 2415 0.945 0.891 0.882 / / / / / / / / / ATRIAL FIBRILLATION 1514 0.993 0.969 0.982 0.925 0.776 0.989 0.953 0.935 0.915 0.950 0.946 0.907 ATRIAL FLUTTER 73 0.993 0.973 0.969 / / / 0.948 0.962 0.937 0.941 0.966 0.956 ELECTRONIC ATRIAL PACEMAKER 298 0.909 0.935 0.745 / / / / / / / / / INFERIOR INFARCT 276 0.851 0.825 0.732 / / / / / / / / / LATERAL INFARCT 1066 0.916 0.819 0.859 / / / / / / / / / LEFT ANTERIOR FASCICULAR BLOCK 167 0.972 0.922 0.919 / / / 0.965 0.882 0.914 0.968 0.819 0.901 LEFT ATRIAL ENLARGEMENT 427 0.799 0.742 0.704 / / / / / / / / / LEFT BUNDLE BRANCH BLOCK 539 0.989 0.969 0.965 0.983 0.963 0.991 0.912 0.932 0.933 0.931 0.917 0.930 LEFT POSTERIOR FASCICULAR BLOCK 18 0.840 0.605 0.904 / / / / / / / / / LEFT VENTRICULAR HYPERTROPHY 2419 0.900 0.860 0.788 / / / / / / / / / LOW VOLTAGE QRS 182 0.581 0.690 0.423 / / / 0.605 0.605 0.814 0.593 0.655 0.741 N-SPECIFIC INTRAVENTRICULAR BLOCK 78 0.766 0.339 0.932 / / / 0.788 0.487 0.906 0.776 0.699 0.793 NORMAL ECG 987 0.887 0.952 0.639 / / / / / / / / / PREMATURE VENTRICULAR COMPLEXES 1146 0.987 0.965 0.959 / / / 0.963 0.955 0.921 0.937 0.939 0.880 QT HAS LENGTHENED 119 0.931 0.897 0.827 / / / 0.930 0.881 0.957 0.930 0.901 0.930 RIGHT ATRIAL ENLARGEMENT 99 0.959 0.869 0.915 / / / / / / / / / RIGHT BUNDLE BRANCH BLOCK 1155 0.976 0.925 0.925 0.967 0.911 0.976 0.932 0.922 0.931 0.946 0.842 0.912 RIGHT VENTRICULAR HYPERTROPHY 136 0.913 0.833 0.859 / / / / / / / / / SEPTAL INFARCT 1423 0.947 0.894 0.880 / / / / / / / / / SINUS BRADYCARDIA 119 0.950 0.972 0.955 0.967 0.938 0.989 0.996 0.966 0.946 0.983 0.935 0.963 SINUS RHYTHM 16785 0.923 0.936 0.751 / / / 0.977 0.942 0.920 0.963 0.926 0.913 SINUS TACHYCARDIA 826 0.994 0.976 0.987 0.988 0.946 0.996 0.986 0.957 0.968 0.979 0.932 0.944 SUPRAVENTRICULAR TACHYCARDIA 27 0.995 0.976 0.959 / / / / / / / / / VENTRICULAR TACHYCARDIA 24 0.987 0.905 0.993 / / / / / / / / / WITH 1ST DEGREE AV BLOCK 82 0.911 0.661 0.921 0.962 0.925 0.992 0.943 0.783 0.911 0.930 0.735 0.893 WITH QRS WIDENING 45 0.618 0.414 0.742 / / / / / / / / / WOLFF-PARKINSON-WHITE 18 0.919 0.747 0.982 / / / / / / / / /
@@
Models ECGFounder Ouyang29
@@ 4 Discussion
ECGFounder enhances diagnostic performance by learning to identify ECG features associated with cardiovascular diseases, which are typically diagnosed based on specific waveform patterns and rhythm characteristics like the elevated ST segments of myocardial infarction and the irregular fluctuations of atrial fibrillation. These features involve anomalies in cardiac electrical activity, appearing significantly different from normal ECG waveforms. Upon training, ECGFounder can recognize these disease-related waveform patterns and rhythms, accurately diagnosing corresponding cardiovascular conditions. As observed in Table S7 in Supplement and Figure 2, ECGFounder matches or even exceeds the performance of cardiology experts on the internal review test sets. Furthermore, ECGFounder’s sensitivity exceeds that of the average cardiology expert, indicating its ability to more accurately capture subtle signs of cardiovascular diseases that may be overlooked by human experts.
@@
Previous research has demonstrated that deep learning models can support clinical ECG analysis and achieve good performance.32, 33, 6 However, existing models lack a universal clinical diagnostic capability for ECGs. Firstly, the training datasets used by existing models are not large or diverse enough, which can lead to overfitting and poor performance on new data, thus limiting their generalizability.34 Additionally, the lack of demographic diversity in the training datasets means these models may perform poorly for certain demographic groups. This could lead to biases when the models are applied to data from different demographic backgrounds, affecting the fairness and accuracy of the models. Secondly, the labels for most model training datasets are derived from manual annotations by cardiology experts, which is time-consuming and labor-intensive. This limits the number of ECGs available for training. Also, because cardiology experts typically use a unified annotation system, the richness of the dataset labels is not very high, often only including common ECG abnormalities and omitting many important but rare diagnostic labels.35 Thirdly, due to the training methods used, these models do not support both 12-lead and single-lead ECGs. To address these issues, we propose a foundation model for ECGs that supports both 12-lead and single-lead ECGs. In Tables 3 and 4, we observed that ECGFounder ranks first in average performance across various external tests. The other baseline models, including S12L-ECG, CTN, and ECG-SE-ResNet, have previously achieved the best performance in AI-ECG models.30, 31, 17 S12L-ECG was trained using supervised learning on the CODE-ECG dataset, which includes 2 million ECGs with 6 common types of arrhythmia diagnoses.17 CTN and ECG-SE-ResNet were trained on the PhysioNet Challenge-2020 dataset, which includes 60,000 ECGs covering 27 common ECG diagnoses.31, 30 In this paper, we demonstrate that by training on a larger and more diverse ECG dataset, a scalable foundation model can be developed to further improve the diagnosis of cardiovascular diseases and surpass previous baseline models.
@@
Despite the effectiveness of ECGFounder in detecting various cardiovascular diseases, there are still challenges to be addressed. Firstly, as most data used to develop ECGFounder come from U.S. cohorts, this limits the diversity and representativeness of the data. Different regions of the world may have unique ECG characteristics, such as race and regional-specific rhythm variations, requiring the model to handle data from diverse backgrounds and populations. Secondly, although ECG foundation models can provide high-accuracy diagnostic results, their decision-making process is often a ’black box,’ which might encounter trust and acceptance issues in clinical applications. Therefore, developing explainable AI models, allowing doctors to understand the model’s decision-making process, is key to advancing the use of ECG AI models. Lastly, some clinically relevant information, such as patients’ lifestyles and medical histories, which could serve as effective covariates in cardiovascular disease research, have not yet been included in the model. We suggest that future work should involve incorporating a larger volume of ECG data from different regions and ethnicities, adding demographic information of patients as model inputs, and developing more explainable AI models, enabling doctors to better understand the decision processes and outcomes of the models. At the same time, some of the latest natural language processing methods have shown great potential in handling ECG annotations from experts, allowing for a more effective use of the clinical knowledge embedded in these annotations for learning.36